OpenStudy (anonymous):

A store mixes Kenyan Coffee worth $10 per kilogram and Turkish coffee worth $14 per kilogram. The mixture is to sell for $11 per kilogram. Find how much of each should be used to make a 204-kilogram mixture.

hero (hero):

x + y = 204 10x + 14y = 11(204)

OpenStudy (anonymous):

so solve for x and y?

hero (hero):

of course

OpenStudy (anonymous):

am i multiplying the whole equation by 204 or just 11?

hero (hero):

you multiply 11*204

OpenStudy (anonymous):

ok that is what i thought.

hero (hero):

11 x 204 = 11*204 = 11(204)

hero (hero):

It's all the same

OpenStudy (anonymous):

2244... ok moving on lol

OpenStudy (anonymous):

so 10x+14y=2244... now i divide or what?

hero (hero):

no, you solve the first equation for y, then substitute

hero (hero):

y = 204 - x

OpenStudy (anonymous):

oh ok...

hero (hero):

?? Did you give up?

OpenStudy (anonymous):

no i was just trying to solve some other problems and get them done

hero (hero):

x + y = 204 10x + 14y = 2244 y = 204 - x 10x + 14(204 - x) = 2244 10x + 2856 - 14x = 2244 -4x + 2856 = 2244 4x = 612 x = 612/4 x = 153 y = 204 - 153 = 51 (153,51)

OpenStudy (anonymous):

GENIUS!!!!!!!!!!!!!!!!!!!!! how do you do it that quick?

hero (hero):

It's basic algebra

OpenStudy (anonymous):

sorry i am just really terrible at math it is just not my subject but i am amazing at english on the other hand

hero (hero):

So am I

OpenStudy (anonymous):

Amazing at english or terrible at math.

hero (hero):

You should really check out the ixl.com site

hero (hero):

I'm good at math and English